# Improved Dynamics for the Maximum Common Subgraph Problem

### 13 Mar 2024 v1

@@ 2.1 Replicator Dynamics (RD)
  • It was found that it’s better to first look for a stationary point using $W=A$ and then use that point to start replicator dynamics from there with $W=\hat{A}$.
@@ 2.4 Kernelization-AIH Pipeline
Kernelization techniques can be added as a processing step to the AIH algorithm. The aim of this addition is to shrink as much as possible the graph to make the execution of AIH faster. Theoretically this is supported by the fact that, choosing the right kernelization techniques, they can be executed in linear time, while AIH algorithm is more complex as they use replicator dynamics and gradient descent. The kernelization algorithm cannot be the same used for its own experiment, because there most graphs get reduced completely and a solution is offered. The aim is only to remove as much vertexes as possible without losing information on a maximum clique in the graph, not to reach a solution; that is left to the AIH algorithm. For this reason the kernelization technique used is the Linear Time algorithm developed by Chang et al. [10] with a slight adjustment. Chang et al.’s algorithm applies a procedure called by them Inexact-Reduction which is very powerful and enables the algorithm to converge to a solution in most cases. This procedure also introduces most of the inaccuracy of the algorithm because it relies on very wide assumptions. Disabling this procedure and avoiding deleting vertexes found to be part of a maximum clique makes the Linear Time algorithm perfectly suited for our purpose, because it reduces the size of the graph but leaves room for AIH to operate with accuracy.
@@ 3.1 AIH vs RD
Experiments are run on randomly generated Erdős-Rényi graphs of size 20 and different levels of expected connectivity. Every pair of graphs randomly generated by this algorithm is then used to build the corresponding association graph. Heuristics RD and AIH are then run on the association graph. Results in Figure 2 show AIH significantly improves over RD.
@@ 3.3 Kernelization-AIH Pipeline
The data in table II show results of the pipeline being consistent with results of the AIH algorithm alone. This validates the pipeline as a viable strategy since similar results have been reached starting from a smaller graph that retains the same information of the bigger, original graph. Due to the nature of how the experiments were conducted the execution times are not really comparable between the single techniques and the pipeline but as explained in section 2 the theory points to a favorable direction in this sense.

### 24 Mar 2024 v2

@@ 2.1 Replicator Dynamics (RD)
  • Empirical results indicate that it’s better to first look for a stationary point using $W=A$ and then use that point to start replicator dynamics from there with $W=\hat{A}$.
@@ 2.4 Kernelization-AIH Pipeline
Kernelization techniques can be added as a processing step to the AIH algorithm. The aim of this addition is to shrink as much as possible the graph to make the execution of AIH faster. Theoretically this is supported by the fact that, choosing the right kernelization techniques, they can be executed in linear time, while AIH algorithm is more complex as they use replicator dynamics and gradient descent. The kernelization algorithm cannot be the same used for its own experiment, because there most graphs get reduced completely and a solution is offered. The aim is only to remove as much vertexes as possible without losing information on a maximum clique in the graph, not to reach a solution; that is left to the AIH algorithm. For this reason the kernelization technique used is the Linear Time algorithm developed by Chang et al. [10] with a slight adjustment. Chang et al.’s algorithm applies a procedure called by them Inexact-Reduction which is very powerful and enables the algorithm to converge to a solution in most cases. This procedure also introduces most of the inaccuracy of the algorithm because it relies on very wide assumptions. Disabling this procedure and avoiding deleting vertexes found to be part of a maximum clique makes the Linear Time algorithm perfectly suited for our purpose, because it reduces the size of the graph but leaves room for AIH to operate with accuracy. Since the kernelization uses the complement of the association graph it is necessary to convert back the kernel before AIH is applied, this procedure is trivial [Figure 1].
@@ 3.1 AIH vs RD
Experiments are run on randomly generated Erdős-Rényi graphs of size 20 and different levels of expected connectivity. Every pair of graphs randomly generated by this algorithm is then used to build the corresponding association graph. Heuristics RD and AIH are then run to find a clique of the association graph as described in the previous section. Results in Figure 3 show AIH significantly improves over RD.
@@ 3.3 Kernelization-AIH Pipeline
The data in TABLE II and Figure 5 show results of the pipeline being consistent with results of the AIH algorithm alone. The result that achieves the best accuracy between the two algorithms is highlighted in bold. Figure 5 shows similar improvement of kernelization+AIH over kernelization+RD compared to the one observed by pure AIH over RD in Figure 3. This validates our pipeline as a viable strategy since similar results have been reached starting from a smaller graph that retains the same information of the bigger, original graph. Due to the nature of how the experiments were conducted execution times are not really comparable between the single techniques and the pipeline but as explained in section 2 theory points to a favorable direction in this sense.